# TxRay: Agentic Postmortem of Live Blockchain Attacks

### v1

@@ Appendix G DeFiHackLabs Postmortem Outcomes
| Date | Incident | ACT | RC | C1 | C2 | C3 | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | Date | Incident | ACT | RC | C1 | C2 | C3 | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | ||||||||||||||||||
| TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | ||||||||
| 24-10 | FireToken | 25-03 | wKeyDAO | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | LavaLending | 25-03 | DCFToken | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | AIZPTToken | 25-03 | BBXToken | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | HYDT | 25-03 | Alkimiya_io | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | SASHAToken | 25-03 | LeverageSIR | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-10 | P719Token | – | – | – | – | – | – | – | – | – | – | 25-04 | AIRWA | ||||||||||||||||||||||||||||||
| 24-10 | MorphoBlue | 25-04 | Laundromat | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-10 | Erc20transfer | 25-04 | Unverified_6077 | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | VISTA | 25-04 | YBToken | – | – | – | – | – | – | – | – | – | – | ||||||||||||||||||||||||||||||
| 24-10 | CompoundFork | 25-04 | BTNFT | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | RPP | 25-04 | Lifeprotocol | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | ChiSale | 25-04 | ImpermaxV3 | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | CoW | 25-05 | Nalakuvara | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | X319 | 25-05 | MBUToken | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | DeltaPrime | – | – | – | – | – | – | – | – | – | – | 25-05 | Unwarp | ||||||||||||||||||||||||||||||
| 24-11 | vETH | 25-05 | KRCToken_pair | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | MFT | 25-05 | RICE | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | MainnetSettler | 25-05 | UsualMoney | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | Matez | 25-05 | Corkprotocol | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-11 | Ak1111 | 25-06 | unverified_8490 | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | proxy_b7e1 | 25-06 | AAVEBoost | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | NFTG | 25-06 | MetaPool | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Pledge | 25-06 | BankrollStack | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | LABUBU | 25-06 | BankrollNetwork | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | CloberDEX | 25-06 | Gangsterfinance | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | JHY | 25-06 | GradientMakerPool | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | BTC24H | 25-06 | unverified_b5cb | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | SlurpyCoin | 25-06 | ResupplyFi | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Moonhacker | 25-06 | Stead | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Bizness | 25-07 | FPC | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | LAURAToken | 25-07 | unverified_54cd | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | sorraStaking | 25-07 | RANTToken | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | 98Token | 25-07 | gmx | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Mosca | 25-07 | VDS | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | IPC | 25-07 | WETC_Token | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | HORS | 25-07 | Stepp2p | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | LPMine | 25-07 | SWAPPStaking | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | JPulsepot | 25-07 | MulticallWithETH | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | RoulettePotV2 | 25-07 | SuperRare | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Unilend | 25-08 | WXC_Token | – | – | – | – | – | – | – | – | – | – | ||||||||||||||||||||||||||||||
| 25-01 | Mosca2 | 25-08 | Bebop_dex | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | IdolsNFT | 25-08 | YuliAI | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Paribus | 25-08 | coinbase | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Ast | 25-08 | Grizzifi | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | ODOS | 25-08 | PDZ | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | PeapodsFinance | 25-08 | SizeCredit | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | FourMeme | 25-08 | d3xai | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | unverified_d4f1 | 25-08 | MulticallWithXera | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | StepHeroNFTs | 25-08 | 0x8d2e | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | unverified_35bc | 25-08 | ABCCApp | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | HegicOptions | 25-08 | 0xf340 | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | Pump | 25-08 | EverValueCoin | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | OneInchFusionV1 | – | – | – | – | – | – | – | – | – | – | 25-08 | Hexotic | ||||||||||||||||||||||||||||||
| 25-03 | UNI | 25-09 | Kame | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | SBRToken | 25-10 | MIMSpell3 | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | DUCKVADER | 25-10 | TokenHolder | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | H2O | 25-11 | BalancerV2 | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| Total | 114 | 109 | 105 | 105 | 114 | 105 | 112 | 105 | 114 | 104 | 109 | 103 | 82 | 53 | 14 | 105 | 3 | 93 | 67 | 105 | 103 | ||||||||||||||||||||||

### v3

@@ Appendix G DeFiHackLabs Postmortem Outcomes
| Date | Incident | ACT | RC | C1 | C2 | C3 | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | |||||||||
| TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | ||||
| 24-10 | FireToken | ||||||||||||||||||||
| 24-10 | LavaLending | ||||||||||||||||||||
| 24-10 | AIZPTToken | ||||||||||||||||||||
| 24-10 | HYDT | ||||||||||||||||||||
| 24-10 | SASHAToken | ||||||||||||||||||||
| 24-10 | P719Token | – | – | – | – | – | – | – | – | – | – | ||||||||||
| 24-10 | MorphoBlue | ||||||||||||||||||||
| 24-10 | Erc20transfer | ||||||||||||||||||||
| 24-10 | VISTA | ||||||||||||||||||||
| 24-10 | CompoundFork | ||||||||||||||||||||
| 24-11 | RPP | ||||||||||||||||||||
| 24-11 | ChiSale | ||||||||||||||||||||
| 24-11 | CoW | ||||||||||||||||||||
| 24-11 | X319 | ||||||||||||||||||||
| 24-11 | DeltaPrime | – | – | – | – | – | – | – | – | – | – | ||||||||||
| 24-11 | vETH | ||||||||||||||||||||
| 24-11 | MFT | ||||||||||||||||||||
| 24-11 | MainnetSettler | ||||||||||||||||||||
| 24-11 | Matez | ||||||||||||||||||||
| 24-11 | Ak1111 | ||||||||||||||||||||
| 24-11 | proxy_b7e1 | ||||||||||||||||||||
| 24-11 | NFTG | ||||||||||||||||||||
| 24-12 | Pledge | ||||||||||||||||||||
| 24-12 | LABUBU | ||||||||||||||||||||
| 24-12 | CloberDEX | ||||||||||||||||||||
| 24-12 | JHY | ||||||||||||||||||||
| 24-12 | BTC24H | ||||||||||||||||||||
| 24-12 | SlurpyCoin | ||||||||||||||||||||
| 24-12 | Moonhacker | ||||||||||||||||||||
| 24-12 | Bizness | ||||||||||||||||||||
| 25-01 | LAURAToken | ||||||||||||||||||||
| 25-01 | sorraStaking | ||||||||||||||||||||
| 25-01 | 98Token | ||||||||||||||||||||
| 25-01 | Mosca | ||||||||||||||||||||
| 25-01 | IPC | ||||||||||||||||||||
| 25-01 | HORS | ||||||||||||||||||||
| 25-01 | LPMine | ||||||||||||||||||||
| 25-01 | JPulsepot | ||||||||||||||||||||
| 25-01 | RoulettePotV2 | ||||||||||||||||||||
| 25-01 | Unilend | ||||||||||||||||||||
| 25-01 | Mosca2 | ||||||||||||||||||||
| 25-01 | IdolsNFT | ||||||||||||||||||||
| 25-01 | Paribus | ||||||||||||||||||||
| 25-01 | Ast | ||||||||||||||||||||
| 25-01 | ODOS | ||||||||||||||||||||
| 25-02 | PeapodsFinance | ||||||||||||||||||||
| 25-02 | FourMeme | ||||||||||||||||||||
| 25-02 | unverified_d4f1 | ||||||||||||||||||||
| 25-02 | StepHeroNFTs | ||||||||||||||||||||
| 25-02 | unverified_35bc | ||||||||||||||||||||
| 25-02 | HegicOptions | ||||||||||||||||||||
| 25-03 | Pump | ||||||||||||||||||||
| 25-03 | OneInchFusionV1 | – | – | – | – | – | – | – | – | – | – | ||||||||||
| 25-03 | UNI | ||||||||||||||||||||
| 25-03 | SBRToken | ||||||||||||||||||||
| 25-03 | DUCKVADER | ||||||||||||||||||||
| 25-03 | H2O | ||||||||||||||||||||
| 25-03 | wKeyDAO | ||||||||||||||||||||
| 25-03 | DCFToken | ||||||||||||||||||||
| 25-03 | BBXToken | ||||||||||||||||||||
| 25-03 | Alkimiya_io | ||||||||||||||||||||
| 25-03 | LeverageSIR | – | – | – | – | – | – | – | – | – | |||||||||||
| 25-04 | AIRWA | ||||||||||||||||||||
| 25-04 | Laundromat | – | – | – | – | – | – | – | – | – | |||||||||||
| 25-04 | Unverified_6077 | ||||||||||||||||||||
| 25-04 | YBToken | – | – | – | – | – | – | – | – | – | – | ||||||||||
| 25-04 | BTNFT | ||||||||||||||||||||
| 25-04 | Lifeprotocol | ||||||||||||||||||||
| 25-04 | ImpermaxV3 | ||||||||||||||||||||
| 25-05 | Nalakuvara | ||||||||||||||||||||
| 25-05 | MBUToken | ||||||||||||||||||||
| 25-05 | Unwarp | ||||||||||||||||||||
| 25-05 | KRCToken_pair | ||||||||||||||||||||
| 25-05 | RICE | ||||||||||||||||||||
| 25-05 | UsualMoney | ||||||||||||||||||||
| 25-05 | Corkprotocol | – | – | – | – | – | – | – | – | – | |||||||||||
| 25-06 | unverified_8490 | ||||||||||||||||||||
| 25-06 | AAVEBoost | ||||||||||||||||||||
| 25-06 | MetaPool | ||||||||||||||||||||
| 25-06 | BankrollStack | ||||||||||||||||||||
| 25-06 | BankrollNetwork | ||||||||||||||||||||
| 25-06 | Gangsterfinance | ||||||||||||||||||||
| 25-06 | GradientMakerPool | ||||||||||||||||||||
| 25-06 | unverified_b5cb | ||||||||||||||||||||
| 25-06 | ResupplyFi | ||||||||||||||||||||
| 25-06 | Stead | ||||||||||||||||||||
| 25-07 | FPC | ||||||||||||||||||||
| 25-07 | unverified_54cd | ||||||||||||||||||||
| 25-07 | RANTToken | ||||||||||||||||||||
| 25-07 | gmx | ||||||||||||||||||||
| 25-07 | VDS | ||||||||||||||||||||
| 25-07 | WETC_Token | ||||||||||||||||||||
| 25-07 | Stepp2p | ||||||||||||||||||||
| 25-07 | SWAPPStaking | ||||||||||||||||||||
| 25-07 | MulticallWithETH | ||||||||||||||||||||
| 25-07 | SuperRare | ||||||||||||||||||||
| 25-08 | WXC_Token | – | – | – | – | – | – | – | – | – | – | ||||||||||
| 25-08 | Bebop_dex | ||||||||||||||||||||
| 25-08 | YuliAI | ||||||||||||||||||||
| 25-08 | coinbase | ||||||||||||||||||||
| 25-08 | Grizzifi | ||||||||||||||||||||
| 25-08 | PDZ | ||||||||||||||||||||
| 25-08 | SizeCredit | ||||||||||||||||||||
| 25-08 | d3xai | ||||||||||||||||||||
| 25-08 | MulticallWithXera | ||||||||||||||||||||
| 25-08 | 0x8d2e | ||||||||||||||||||||
| 25-08 | ABCCApp | ||||||||||||||||||||
| 25-08 | 0xf340 | ||||||||||||||||||||
| 25-08 | EverValueCoin | ||||||||||||||||||||
| 25-08 | Hexotic | ||||||||||||||||||||
| 25-09 | Kame | ||||||||||||||||||||
| 25-10 | MIMSpell3 | ||||||||||||||||||||
| 25-10 | TokenHolder | ||||||||||||||||||||
| 25-11 | BalancerV2 | – | – | – | – | – | – | – | – | – | |||||||||||
| Total | 114 | 109 | 105 | 105 | 114 | 105 | 112 | 105 | 114 | 104 | 112 | 103 | 85 | 53 | 17 | 105 | 3 | 93 | 67 | 105 | 103 |